OpenStudy (trisarahtops):

FAN and MEDAL! If h(x) = f[f(x)] use the table of values for f and f ′ to find the value of h ′(1). Here is the chart given x f(x) f ′(x) 1 3 2 2 1 5 3 6 7

OpenStudy (anonymous):

\[\left(f(f)\right)'=f'(f)\times f'\] by the chain rule you have all the numbers you need to find \[f'(f(1))\times f'(1)\]

OpenStudy (trisarahtops):

2(3)*2?

OpenStudy (trisarahtops):

@satellite73

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

ok actually i do not have to check to know it is wrong

OpenStudy (anonymous):

the first part is a composition of functions \[f'(f(1))\]

OpenStudy (anonymous):

what is \(f(1)\)?

OpenStudy (trisarahtops):

3

OpenStudy (anonymous):

ok so \[f'(f(1))=f'(3)=?\]

OpenStudy (trisarahtops):

so you saying it's not like f'(1)* f(1)

OpenStudy (anonymous):

no there is a product there, but also a composition of functions

OpenStudy (anonymous):

\[f'(f(1))\times f'(1)\]

OpenStudy (anonymous):

the \(\times \) is a multiplication but \[f'(f(1))\] is a composition of functions "f prime of f of 1"

OpenStudy (trisarahtops):

sorry I took so long I left my computer for a bit. f'(f1) = f'(3)= 7? @satellite73

OpenStudy (trisarahtops):

sorry I meant f'(f(1))= f'(3)= 7

OpenStudy (trisarahtops):

@satellite73 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then multiply by \(f'(1)\) to finish

OpenStudy (trisarahtops):

7*2 = 14

OpenStudy (trisarahtops):

so 14 is the final anwser correct? @satellite73